Given the following information regarding direct materials:   ---Actual quantity purchased = 75,000 pounds -Actual quantity used = 66,000 pounds   Actual price of direct materials = $11 per pound Standard quantity of direct materials = 3 pounds per unit Standard cost of direct materials = $10 per pound Units produced = 21,000 units

-What is the direct materials price variance? (AP - SP) x AQ purchased ?

-What is the direct materials usage variance? (AQ used - SQ allowed) x SP ?
Answer #1

(a) Direct material price variance = (Actual price - Standard price) * Actual quantity

Standard price = $10 per pound

Actual price = $11 per pound

Actual quantity = 75000 pounds

Now, putting these values in the above formula, we get,

Direct material price variance = ($11 - $10) * 75000

Direct material price variance = $1 * 75000 = $75000

It is an unfavorable variance.

(b) Direct material usage variance = (Actual quantity - Standard quantity allowed) * Standard price

Standard price = $10 per pound

Standard quantity allowed = Units produced * Standard quantity per unit

Standard quantity allowed = 21000 * 3 = 63000

Actual quantity (used) = 66000 pounds

Now, putting these values in the above formula, we get,

Direct material usage variance = (66000 - 63000) * $10

Direct material usage variance = 3000 * $10 = $30000

It is an unfavorable variance.